OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12

OpenStudy (whpalmer4):

\[-x-2y=-4\]You want x all by itself, and you want it to be +x, not -x, so I would suggest adding x to both sides, and adding 4 to both sides.

OpenStudy (whpalmer4):

you'll get an equation with x on the right side instead of left, but because they are equal, you can reverse the equation.

OpenStudy (anonymous):

2y = -4 + x ?

OpenStudy (whpalmer4):

No. \[-x-2y-4\]Add x to both sides of that, what do you get? No need to tag me, if I've ever responded to the question, I'll get a notification when someone does anything.

OpenStudy (whpalmer4):

ah, \[-x-2y= -4\]

OpenStudy (anonymous):

I dont know? brain isnt funtioning right today.. Sorry

OpenStudy (whpalmer4):

Come on. \[-x-2y=-4\]Add \(x\) to each side: \[-x - 2y + x = -4 + x\]simplify that

OpenStudy (whpalmer4):

You do that.